# Wigner’s Theorem for stabilizer states and quantum designs

### v1

@@ I.2.1 The Pauli group
$(A|B):=\operatorname{tr}A^{\dagger}B,$
where we used the notation adopted in scott2006tight . Using Eq. (2), one may easily verify that the Weyl-Heisenberg operators are ortho-normal with respect to this inner product:
$\big{(}T(\mathbb{a})\big{|}T(\mathbb{b})\big{)}=d^{n}\delta_{\mathbb{a},\mathbb{b}}.$ (4)
@@ II.1 The symmetries of the stabilizer polytope
The convex hull of the projectors of a set of $d+1$ mutually unbiased bases is referred to as the complementarity polytope in the literature Bengtsson2006 ; Bengtsson2004 ; Appleby2015GaloisUM ; dang2015studies . The fact that the symmetry group of the complementarity polytope has the structure (as a set) of $S_{d+1}\times S_{d}\times\dots\times S_{d}$ has been stated in the papers just cited. The group structure $S_{d}\wr S_{d+1}$ has been identified before as the combinatorial symmetry group of $d+1$ mutually unbiased bases Gribling2021 .
@@ III.1 Moments of $Q$
$F_{k}^{Q}(A_{1},\dots,A_{k}):=\big{(}A_{1}\otimes\dots\otimes A_{k}\big{|}\mu_{k}^{Q}\big{)}=\int_{Q}\big{(}A_{1}\big{|}q\big{)}\dots\big{(}A_{k}\big{|}q\big{)}\,\mathrm{d}\nu(q).$ (20)
@@
$F^{Q}_{k}(\Lambda^{\dagger}(A_{1}),\dots,\Lambda^{\dagger}(A_{k}))=$ $\int\big{(}\Lambda^{\dagger}(A_{1})\big{|}q\big{)}\dots\big{(}\Lambda^{\dagger}(A_{k})\big{|}q\big{)}\,\mathrm{d}\nu(q)$
$=$ $\int\big{(}A_{1}\big{|}\Lambda(q)\big{)}\dots\big{(}A_{k}\big{|}\Lambda(q)\big{)}\,\mathrm{d}\nu(q)=F_{k}(A_{1},\dots,A_{k}).$
@@
$F_{2}^{Q}(A,B)=\frac{1}{d(d+1)}\big{(}\operatorname{tr}A\,\operatorname{tr}B+\operatorname{tr}AB\big{)}=\frac{1}{d(d+1)}\big{(}(A|B)+(A|\mathbbm{1})(\mathbbm{1}|B)\big{)}.$ (23)
@@
$F_{3}^{Q}(A,B,C)=\frac{1}{\dim(\operatorname{Sym}^{3}(\mathcal{H}))}\Big{(}(A|\mathbbm{1})(\mathbbm{1}|B)(\mathbbm{1}|C)+(A|\mathbbm{1})(B|C)+(B|\mathbbm{1})(A|C)+(C|\mathbbm{1})(A|B)+\operatorname{tr}(ABC)+\operatorname{tr}(ACB)\Big{)}.$
@@
$F^{Q}_{2}(A_{1},A_{2})=\frac{1}{d^{2}}\sum_{\mathbb{a}}\big{(}A_{1}|A(\mathbb{a})\big{)}\big{(}A(\mathbb{a})|A_{2}\big{)}=\frac{1}{d^{2}}\big{(}A_{1}|A_{2}\big{)}.$
@@ III.1.1 Real spherical designs
$F_{3}^{Q}(A,B,C)=K_{1}(A|\mathbbm{1})(\mathbbm{1}|B)(\mathbbm{1}|C)+K_{2}\Big{(}(A|\mathbbm{1})(B|C)+(B|\mathbbm{1})(A|C)+(C|\mathbbm{1})(A|B)\Big{)}+K_{3}\Big{(}\operatorname{tr}(ABC)+\operatorname{tr}(ACB)\Big{)}$ (25)
@@ III.2 Relations between symmetry groups
The following statement seems to be commonly known. We re-prove it because we do not know of a suitable self-contained presentation in the literature.
@@
$\big{(}q_{i}\big{|}W(q)\big{)}=\big{(}W^{-1}(q_{i})\big{|}q\big{)}=\sum_{j}c_{j}\big{(}W^{-1}(q_{i})\big{|}q_{j}\big{)}=\sum_{j}c_{j}\operatorname{tr}\big{(}q_{i}\big{|}W(q_{j})\big{)}=\big{(}q_{i}|L(q)\big{)}.$
@@ III.2.1 Implications based on moments
$\begin{split}&F^{Q}_{2}\text{ is proportional to }(\cdot|\cdot)\text{ when restricted to }\operatorname{dir}(Q)\\ \text{and }\qquad&\mu^{Q}_{1}\text{ is orthogonal to $\operatorname{dir}(Q)$ with respect to both $F^{Q}_{2}$ and the Hilbert-Schmidt form}\end{split}$ (27)
@@ IV.1.1 Proof via Wigner symmetries
$\left(\Pi_{L}^{g}\big{|}\Pi_{M}^{h}\right)=\begin{cases}1,&\text{if }L=M\text{and }g=h\\ 0,&\text{if }L=M\text{ and }g\neq h\\ \frac{1}{d}&\text{if }L\neq M,\end{cases}$ (31)
@@ IV.3 $d$ odd, $n>1$
$\left(\Pi_{L}^{g}\big{|}\Pi_{M}^{h}\right)=d^{\dim{(L\cap M)}-2n}\ \delta\left[g|_{L\cap M}=h|_{L\cap M}\right].$ (41)
@@ Appendix A Spherical designs in Euclidean spaces
$F^{Q}_{k}(A_{1},\ldots,A_{k})=\int_{Q}\big{(}A_{1}\big{|}q\big{)}\cdots\big{(}A_{k}\big{|}q\big{)}\,\mathrm{d}\nu(q)=\frac{1}{|X|}\sum_{x\in X}A_{1}[x]\cdots A_{k}[x],$
@@
see for example [46, Lemma 3.5.3]  [^1] The factor $2$ in (3.5.11) is not correct in [46]; it should be $1$.. Thus, applying the Laplace-operator $k$-times to (54) yields the following identity:
@@
$F_{2}^{Q}(A,B)=\frac{1}{|X|}\sum_{x\in X}A[x]B[x]=\frac{1}{|X|}\sum_{x\in X}p_{0}(x)=\frac{1}{|X|}\sum_{x\in X}K\Delta^{2}(A[x]B[x])=K^{\prime}\Big{(}(A|B)+2(A|\mathbbm{1})(B|\mathbbm{1})\Big{)}.$

### v2

@@ I.2.1 The Pauli group
$\operatorname{tr}A^{\dagger}B,$
Using Eq. (2), one may easily verify that the Weyl-Heisenberg operators are orthonormal with respect to this inner product:
$\operatorname{tr}T(\mathbb{a})^{\dagger}T(\mathbb{b})=d^{n}\delta_{\mathbb{a},\mathbb{b}}.$ (4)
@@ II.1 The symmetries of the stabilizer polytope
The convex hull of the projectors of a set of $d+1$ mutually unbiased bases is referred to as the complementarity polytope in the literature Bengtsson2006 ; Bengtsson2004 ; Appleby2015GaloisUM ; dang2015studies . The fact that the symmetry group of the complementarity polytope has the structure (as a set) of $S_{d+1}\times S_{d}\times\dots\times S_{d}$ has been stated in the papers just cited. The group structure $S_{d}\wr S_{d+1}$ has been identified before as the combinatorial symmetry group of $d+1$ mutually unbiased bases Gribling2021 . Note that for the $d=2$, $n=1$ case, the symmetry group, the wreath product structure $S_{2}\wr\S_{3}$ is the same as the extended Clifford group (see Appendix B).
@@ III.1 Moments of $Q$
$F_{k}^{Q}(A_{1},\dots,A_{k}):=\operatorname{tr}\big{(}A_{1}\otimes\dots\otimes A_{k}\ \mu_{k}^{Q}\big{)}=\int_{Q}\operatorname{tr}\big{(}A_{1}q\big{)}\dots\operatorname{tr}\big{(}A_{k}q\big{)}\,\mathrm{d}\nu(q).$ (20)
@@
$F^{Q}_{k}(\Lambda^{\dagger}(A_{1}),\dots,\Lambda^{\dagger}(A_{k}))=$ $\int\operatorname{tr}\big{(}\Lambda^{\dagger}(A_{1})\ q\big{)}\dots\operatorname{tr}\big{(}\Lambda^{\dagger}(A_{k})\ q\big{)}\,\mathrm{d}\nu(q)$
$=$ $\int\operatorname{tr}\big{(}A_{1}\ \Lambda(q)\big{)}\dots\operatorname{tr}\big{(}A_{k}\ \Lambda(q)\big{)}\,\mathrm{d}\nu(q)=F_{k}(A_{1},\dots,A_{k}).$
@@
$F_{2}^{Q}(A,B)=\frac{1}{d(d+1)}\big{(}\operatorname{tr}A\,\operatorname{tr}B+\operatorname{tr}AB\big{)}$ (23)
@@
$F_{3}^{Q}(A,B,C)=\frac{1}{\dim(\operatorname{Sym}^{3}(\mathcal{H}))}\Big{(}\operatorname{tr}A\operatorname{tr}B\operatorname{tr}C+\operatorname{tr}A\operatorname{tr}BC+\operatorname{tr}B\operatorname{tr}AC+\operatorname{tr}C\operatorname{tr}AB+\operatorname{tr}(ABC)+\operatorname{tr}(ACB)\Big{)}.$
@@
$F^{Q}_{2}(A_{1},A_{2})=\frac{1}{d^{2}}\sum_{\mathbb{a}}\operatorname{tr}\left(A_{1}A(\mathbb{a})\right)\operatorname{tr}\left(A(\mathbb{a})A_{2}\right)=\frac{1}{d^{2}}\operatorname{tr}A_{1}A_{2}.$
@@ III.1.1 Real spherical designs
$F_{3}^{Q}(A,B,C)=K_{1}\operatorname{tr}A\operatorname{tr}B\operatorname{tr}C+K_{2}\Big{(}\operatorname{tr}A\operatorname{tr}BC+\operatorname{tr}B\operatorname{tr}AC+\operatorname{tr}C\operatorname{tr}AB\Big{)}+K_{3}\Big{(}\operatorname{tr}(ABC)+\operatorname{tr}(ACB)\Big{)}$ (25)
@@ III.2 Relations between symmetry groups
The following statement seems to be commonly known. A related argument appears e.g. in the proof of Prop. 5.15 in Ref. landsman2017 . We re-prove it because we have not found a suitable self-contained presentation in the literature.
@@
The next claim is essentially a special case of the Mazur-Ulam Theorem.
@@
$\operatorname{tr}\big{(}q_{i}\ W(q)\big{)}=\operatorname{tr}\big{(}W^{-1}(q_{i})\ q\big{)}=\sum_{j}c_{j}\operatorname{tr}\big{(}W^{-1}(q_{i})\ q_{j}\big{)}=\sum_{j}c_{j}\operatorname{tr}\big{(}q_{i}\ W(q_{j})\big{)}=\operatorname{tr}\big{(}q_{i}\ L(q)\big{)}.$
@@ III.2.1 Implications based on moments
$\begin{split}&F^{Q}_{2}\text{ is proportional to the Hilbert-Schmidt form when restricted to }\operatorname{dir}(Q)\\ \text{and }\qquad&\mu^{Q}_{1}\text{ is orthogonal to $\operatorname{dir}(Q)$ with respect to both $F^{Q}_{2}$ and the Hilbert-Schmidt form}\end{split}$ (27)
@@ IV.1.1 Proof via Wigner symmetries
$\operatorname{tr}\left(\Pi_{L}^{g}\ \Pi_{M}^{h}\right)=\begin{cases}1,&\text{if }L=M\text{ and }g=h\\ 0,&\text{if }L=M\text{ and }g\neq h\\ \frac{1}{d}&\text{if }L\neq M,\end{cases}$ (31)
@@ IV.3 $d$ odd, $n>1$
$\operatorname{tr}\left(\Pi_{L}^{g}\ \Pi_{M}^{h}\right)=d^{\dim{(L\cap M)}-2n}\ \delta\left[g|_{L\cap M}=h|_{L\cap M}\right].$ (41)
@@ Appendix A Spherical designs in Euclidean spaces
$F^{Q}_{k}(A_{1},\ldots,A_{k})=\int_{Q}\operatorname{tr}\big{(}A_{1}q\big{)}\cdots\operatorname{tr}\big{(}A_{k}q\big{)}\,\mathrm{d}\nu(q)=\frac{1}{|X|}\sum_{x\in X}A_{1}[x]\cdots A_{k}[x],$
@@
see for example [45, Lemma 3.5.3]   [^1] The factor $2$ in (3.5.11) is not correct in [45]; it should be $1$ Thus, applying the Laplace-operator $k$-times to (54) yields the following identity:
@@
$F_{2}^{Q}(A,B)=\frac{1}{|X|}\sum_{x\in X}A[x]B[x]=\frac{1}{|X|}\sum_{x\in X}p_{0}(x)=\frac{1}{|X|}\sum_{x\in X}K\Delta^{2}(A[x]B[x])=K^{\prime}\Big{(}\operatorname{tr}AB+2\operatorname{tr}A\operatorname{tr}B\Big{)}.$